### v1

# Language Models as Hierarchy Encoders

Abstract
Interpreting hierarchical structures latent in language is a key limitation of current language models (LMs). While previous research has implicitly leveraged these hierarchies to enhance LMs, approaches for their explicit encoding are yet to be explored. To address this, we introduce a novel approach to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs), harnessing the expansive nature of hyperbolic space. Our method situates the output embedding space of pre-trained LMs within a Poincaré ball with a curvature that adapts to the embedding dimension, followed by re-training on hyperbolic cluster and centripetal losses. These losses are designed to effectively cluster related entities (input as texts) and organise them hierarchically. We evaluate HiTs against pre-trained and fine-tuned LMs, focusing on their capabilities in simulating transitive inference, predicting subsumptions, and transferring knowledge across hierarchies. The results demonstrate that HiTs consistently outperform both pre-trained and fine-tuned LMs in these tasks, underscoring the effectiveness and transferability of our re-trained hierarchy encoders.222 Code on GitHub: https://github.com/KRR-Oxford/HierarchyTransformers; Datasets on Zenodo: https://zenodo.org/doi/10.5281/zenodo.10511042; HiT models on Huggingface Hub: https://huggingface.co/Hierarchy-Transformers
Language Models, Natural Language Processing, Deep Learning, Machine Learning, ICML
1 Introduction
Figure 1: Illustration of how hierarchies are explicitly encoded in HiT after re-training. The square ($d$-dimensional hyper-cube) refers to the output embedding space of transformer encoder-based LMs whose final activation function is $\tanh$, and the circumscribed circle ($d$-dimensional hyper-sphere) refers to the Poincaré ball of radius $\sqrt{d}$. The distance and norm metrics involved in our re-training hyperbolic losses are defined w.r.t. this manifold.
In the field of Natural Language Processing (NLP) and related areas, the emergence of transformer-based language models (LMs) such as BERT (encoder-based) (Kenton & Toutanova, 2019), GPT (decoder-based) (Radford et al., 2018), and the more recent large language models (LLMs) like GPT-4 (OpenAI, 2023) and Llama 2 (Touvron et al., 2023), has marked a significant progression. Nonetheless, these models face a notable challenge in effectively encoding and interpreting hierarchical structures latent in language. This limitation has been highlighted by several studies, including those by Hanna & Mareček (2021) and He et al. (2023b), which employed prompt-based probes to reveal the limited hierarchical knowledge in pre-trained LMs, and the work by Lin & Ng (2022), which demonstrated these models’ struggles with capturing the transitivity of hierarchical relationships.
Prior research has explored various methods to infuse hierarchical information into LM training. Common approaches include classification-based fine-tuning using sentence head embedding with a classification layer (Chen et al., 2023) or few-shot prompting with an answer mapping to classification labels (He et al., 2023b). To pre-train, or re-train [^3] The term re-train refers to further pre-train LMs on another corpus; it is distinguished from fine-tune as it does not modify the LM’s architecture, while the latter often involves adding task-specific layers which lead to additional learnable parameters. LMs on a corpus constructed from hierarchical data, Liu et al. (2020) converted structural representations into textual formats to align the masked language modeling objective. Others, like Liu et al. (2021) and Gosselin & Zouaq (2023), have focused on extracting analogous and contrasting examples from hierarchical structures for a similarity-based contrastive learning objective. The aforementioned studies leveraged hierarchical information as implicit signals to augment LMs, yet no existing works specifically targeted the explicit encoding of hierarchies with LMs.
To bridge this gap, we introduce a novel approach to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs). Inspired by the efficacy of hyperbolic geometry in representing hierarchical structures (Nickel & Kiela, 2017; Ganea et al., 2018b), we propose the hyperbolic clustering and centripetal losses tailored for LM re-training. As illustrated in Figure 1, transformer encoder-based LMs typically use a $\tanh$ activation function in the last layer, which maps each embedding dimension to the range $[-1,1]$. Consequently, the output embeddings of LMs are confined within a unit $d$-dimensional hyper-cube. Leveraging this characteristic, we utilise a Poincaré ball of radius $\sqrt{d}$, whose boundary circumscribes [^4] We ignore the vertices of the hyper-cube, as they are on the boundary and thus are undefined in the open Poincaré ball. the output embedding space of LMs. The metrics for distance and norm used in our hyperbolic losses are defined w.r.t. this specific manifold. After re-training, entities are not only clustered according to their relatedness but also hierarchically organised.
In evaluating HiTs, we compare their performance against both pre-trained and fine-tuned LMs in the Multi-hop Inference and Mixed-hop Prediction tasks. The Multi-hop Inference task, following the setting in Ganea et al. (2018b), involves training models on all asserted (i.e., one-hop) subsumptions and assessing their ability to infer transitive (i.e., multi-hop) subsumptions. The Mixed-hop Prediction task is designed to mirror real-world scenarios, where models trained on incomplete hierarchy are applied to predict unknown subsumption relationships between arbitrary entity pairs. Additionally, we introduce a transfer learning setting, where models trained on one hierarchy are tested on another. Our experiments utilise datasets derived from WordNet’s noun hierarchy (Miller, 1995) and the SNOMED CT ontology (Stearns et al., 2001), [^5] Results on WordNet (Noun) are presented in Section 4, while results on SNOMED CT are presented in Appendix C. and transfer evaluation datasets from Schema.org (Guha et al., 2016), Food Ontology (FoodOn) (Dooley et al., 2018), and Disease Ontology (DOID) (Schriml et al., 2012). The results show that HiTs significantly surpass both pre-trained and fine-tuned LMs in these tasks, demonstrating their robustness to generalise from asserted to inferred and unseen subsumptions, and a promising potential in hierarchy-oriented semantic search.
2 Preliminaries
Before delving into our methodology, we provide a brief overview of key concepts: transformer encoder-based LMs, hyperbolic geometry, and the formal definition of hierarchy.
2.1 Language Models
Transformer encoder-based LMs excel in providing fine-grained contextual word embeddings for enhanced language understanding. A key component of these models is the self-attention mechanism, which dynamically assigns importance to different segments of the input text, thereby capturing nuanced contextual semantics more effectively. Notable examples of such models include BERT (Kenton & Toutanova, 2019) and RoBERTa (Liu et al., 2019), both of which utilise the masked language modelling objective during pre-training. This approach involves partially masking input sentences and prompting the model to predict the masked tokens, using the unmasked surrounding text as context. For acquiring sentence-level embeddings, these models can be augmented with an additional pooling layer, applied over the token embeddings (Xiao, 2018; Reimers & Gurevych, 2019). Pooling strategies such as mean, max, and sentence head pooling are employed, with their effectiveness varying across different applications. A contrastive learning objective is often applied for refining sentence-level semantics (Reimers & Gurevych, 2019; Gao et al., 2021). Despite the rise of generative LLMs, transformer encoder-based LMs maintain their importance, offering versatility and efficiency in tasks like text classification and semantic search.
2.2 Hyperbolic Geometry
Hyperbolic geometry, a form of non-Euclidean geometry, is characterised by its constant negative Gaussian curvature, a fundamental aspect that differentiates it from the flat, zero curvature of Euclidean geometry. In hyperbolic space, distances between points increase exponentially as one moves towards the boundary, making it inherently suitable for embedding hierarchical structures. This intuition aligns with the tree embedding theorem based on $\delta$-hyperbolicity, as discussed in Gromov (1987) and Ganea et al. (2018b).
Among the various models [^6] E.g., the Poincaré ball model, the Poincaré half plane model, and the hyperboloid model. of hyperbolic geometry that are isometric [^7] An isometry is a bijective distance-preserving transformation. to each other, the Poincaré ball is chosen for its capacity to contain the the output embedding space of LMs directly, as explained in the second last paragraph of Section 1. The $d$-dimensional Poincaré ball with a negative curvature $-c$ (where $c>0$) is defined by the open ball $\mathbb{B}_{c}^{d}=\{\mathbf{x}\in\mathbb{R}^{d}:\lVert\mathbf{x}\rVert^{2}<\frac{1}{c}\}$. The distance function in this model, dependent on the curvature value $c$, is given by:
$d_{c}(\mathbf{u},\mathbf{v})=\frac{2}{\sqrt{c}}\tanh^{-1}(\sqrt{c}\lVert-\mathbf{u}\oplus_{c}\mathbf{v}\rVert)$ (1)
[…]
$\mathbf{u}\oplus_{c}\mathbf{v}=\frac{(1+2c\langle\mathbf{u},\mathbf{v}\rangle+c\lVert\mathbf{v}\rVert^{2})\mathbf{u}+(1-c\lVert\mathbf{u}\rVert^{2})\mathbf{v}}{1+2c\langle\mathbf{u},\mathbf{v}\rangle+c^{2}\lVert\mathbf{u}\rVert^{2}\lVert\mathbf{v}\rVert^{2}}$ (2)
[…]
2.3 Hierarchy
We define a hierarchy as a directed acyclic graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ where $\mathcal{V}$ represents vertices that symbolise entities, and $\mathcal{E}$ represents edges that indicate the direct subsumption relationships asserted in the hierarchy. We can then derive indirect subsumption relationships $\mathcal{T}$ based on direct ones through transitive reasoning. We borrow the notation from description logic to denote the subsumption relationship as $e_{1}\sqsubseteq e_{2}$, meaning that $e_{1}$ is a sub-class of $e_{2}$. Under the closed-world assumption, we consider an edge $(e_{1},e_{2})$ as a negative sample if $(e_{1},e_{2})\not\in\mathcal{E}\cup\mathcal{T}$. Particularly, a hard negative is identified when $e_{1}$ and $e_{2}$ are also siblings that share the same parent.
Explicit hierarchies can often be derived from structured data sources such as taxonomies, ontologies, and knowledge graphs. A taxonomy and the Terminology Box (TBox) of an ontology intrinsically define subsumptions, whereas in knowledge graphs, hierarchical relationships are defined in a more customised manner. For instance, in WordNet, the hypernym relationship corresponds to subsumption.
3 Hierarchy Transformer Encoder
We intend to propose a general and effective strategy to re-train transformer encoder-based LMs as Hierarchy Transformer encoders (HiTs). To deal with arbitrary input lengths of entity names, we employ an architecture similar to sentence transformers (Reimers & Gurevych, 2019), incorporating a mean pooling layer over token embeddings to produce sentence embeddings for entities. Note that some of the sentence transformer models have a normalisation layer after pooling; we exclude this layer because its presence will constrain the embeddings’ Euclidean norms to one, thus hindering hierarchical organisation of entity embeddings. It is also worth mentioning that these changes do not add in learnable parameters besides the ones already in LMs.
As aforementioned, the output embedding space of these LMs is typically a $d$-dimensional hyper-cube because of the $\tanh$ activation function in the last layer. Thus, we can construct a Poincaré ball of radius $\sqrt{d}$ (or equivalently, curvature value $c=\frac{1}{d}$) whose boundary circumscribes the hyper-cube. [^8] We have also considered scaling down each dimension of the LM embeddings by $\sqrt{d}$ to confine them within a unit Poincaré ball, but we found that losses are harder to converge in this construction. Based on this manifold, we propose the following two losses for hierarchy re-training.
Figure 2: Illustration of the impact of hyperbolic clustering and centripetal losses during training. In Euclidean space, it seems contradictory that both “phone” and “computer” are pulled towards “e-device” but are also pushed away from each other. However, in principle this is not a problem in hyperbolic space, where distances increase exponentially relative to Euclidean distances as one moves from the origin to the boundary of the manifold.
Hyperbolic Clustering Loss
This loss aims at clustering related entities and distancing unrelated ones in the Poincaré ball. We formulate it in the form of triplet loss because related entities are not equivalent but their semantic distances should be smaller than those between unrelated entities. Formally, the loss is defined as:
$\mathcal{L}_{cluster}=\sum_{(e,e^{+},e^{-})\in\mathcal{D}}\max(d_{c}(\mathbf{e},\mathbf{e^{+}})-d_{c}(\mathbf{e},\mathbf{e^{-}})+\alpha,0)$ (3)
Here, inputs are presented in the form of triplet $(e,e^{+},e^{-})$, where $e^{+}$ is a parent entity of $e$, and $e^{-}$ is a negative parent entity of $e$; $\mathcal{D}$ denotes the set of these triplets; $d_{c}(\cdot,\cdot)$ refers to the hyperbolic distance function defined in equation 1, and $\alpha$ is the hyperbolic distance margin. The bold letters denote the embeddings of the corresponding entities.
Hyperbolic Centripetal Loss
This loss ensures parent entities are positioned closer to the Poincaré ball’s origin than their child counterparts, reflecting the natural expansion of hierarchies from the origin to the boundary of the manifold. The term “centripetal” is used to imply that the manifold’s origin represents an imaginary root entity for everything. Formally, the hyperbolic centripetal loss is defined as:
$\mathcal{L}_{centri}=\sum_{(e,e^{+},e^{-})\in\mathcal{D}}\max(\lVert\mathbf{e}^{+}\rVert_{c}-\lVert\mathbf{e}\rVert_{c}+\beta,0)$ (4)
Again, inputs are the triplets in $\mathcal{D}$, but only the child and parent entities (the positive subsumptions) are used to calculate the loss; $\lVert\cdot\rVert_{c}:=d_{c}(\cdot,\mathbf{0})$ refers to the hyperbolic norm; $\beta$ is the hyperbolic norm margin.
The overall hierarchy re-training loss, denoted as $\mathcal{L}_{\textsc{HiT}}$, is the linear combination of these two hyperbolic losses, defined as:
$\mathcal{L}_{\textsc{HiT}}=\mathcal{L}_{cluster}+\mathcal{L}_{centri}$ (5)
In Figure 2, we demonstrate the impact of $\mathcal{L}_{\textsc{HiT}}$ on entity embeddings. The entity “e-device”, being most general, is nearest to the origin. Sibling entities, such as “phone” and “computer”, “laptop” and “pc”, are closer to their common parent than to each other, illustrating the effect of retraining to cluster related entities while maintain hierarchical relationships.
As the HiT model functions as an encoder, it does not inherently support direct predictions of subsumption relationships. To address this, we devise an empirical scoring function that leverages the hierarchy re-trained entity embeddings. This function aims to predict the subsumption relationship for any given pair of entities $(e_{1},e_{2})$, incorporating both the clustering and centripetal heuristics:
$s(e_{1}\sqsubseteq e_{2})=-(d_{c}(\mathbf{e}_{1},\mathbf{e}_{2})+\lambda(\lVert\mathbf{e}_{2}\rVert_{c}-\lVert\mathbf{e}_{1}\rVert_{c}))$ (6)
Here, $\lambda>0$ represents a weighting factor applied to the centripetal heuristic component. The subsumption score is structured to increase as the hyperbolic distance between $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ decreases, and/or as the relative difference in their hyperbolic norms ($\lVert\mathbf{e}_{1}\rVert_{c}-\lVert\mathbf{e}_{2}\rVert_{c}$) increases. Essentially, for the model to predict $e_{1}\sqsubseteq e_{2}$, it is expected that $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ are relatively closer in the Poincaré ball, with $\mathbf{e}_{1}$ positioned further from the manifold’s origin compared to $\mathbf{e}_{2}$. The value for $\lambda$ and the overall scoring threshold are to be ascertained through hyperparameter tuning on the validation set.
4 Evaluation
Source #Entity #DirectSub #IndirectSub #Dataset (Train/Val/Test)
WordNet 74,401 75,850 587,658 multi: 834K///323K///323K
mixed: 751K///365K///365K
Schema.org 903 950 1,978 mixed: -///15K///15K
FoodOn 30,963 36,486 438,266 mixed: 361K///261K///261K
DOID 11,157 11,180 45,383 mixed: 122K///31K///31K
Table 1: Statistics of WordNet (Noun), Schema.org, and FoodOn, including the numbers of entities (#Entity), direct subsumptions (#DirectSub), indirect subsumptions (#IndirectSub), and the dataset splittings (#Dataset) for Multi-hop Inference and Mixed-hop Prediction tasks. Note that the numbers in #Dataset are counts of entity pairs rather than entity triplets.
4.1 Task Definition
In our evaluation, we propose the following two tasks to assess a model’s ability to generalise from asserted to inferred and unseen subsumptions.
Multi-hop Inference
This task, following the setting in Ganea et al. (2018b), aims to evaluate the model’s ability in deducing indirect, multi-hop subsumptions $\mathcal{T}$ from direct, one-hop subsumptions $\mathcal{E}$, so as to simulate transitive inference. We split $\mathcal{T}$ for validation and testing, denoted as $\mathcal{T}_{val}$ and $\mathcal{T}_{test}$, respectively. For each positive subsumption $(e,e^{+})$ involved, we sampled $10$ negative parents $e^{-}$ for $e$, leading to $10$ training triplets99910101010 training triplets are constructed from $11$ entity pairs.. Following the criteria in Section 2.3, $(e,e^{-})$ is a valid negative if $(e,e^{-})\not\in\mathcal{E}\cup\mathcal{T}$. We further split the task into two settings: one with random negatives and another with hard negatives, the latter mainly comprising sibling entities. Since not every entity has enough siblings, we supplemented with random negatives that have been sampled to maintain a consistent positive-to-negative ratio of $1:10$.
Mixed-hop Prediction
This task aims to evaluate the model’s capability in determining the existence of subsumption relationships between arbitrary entity pairs, where the entities are not necessarily seen during training. We propose a challenging setting where models are trained on incomplete direct subsumptions and examined on a mix of hold-out, unseen direct and indirect (mixed-hop) subsumptions. We split $\mathcal{E}$ into training, validation, and test sets, denoted as $\mathcal{E}_{train}$, $\mathcal{E}_{val}$, and $\mathcal{E}_{test}$, respectively. The final training, validataion, and test sets for this task are $\mathcal{E}_{train}$, $\mathcal{E}_{val}\cup\mathcal{T}_{val}$, and $\mathcal{E}_{test}\cup\mathcal{T}_{test}$, respectively, where $\mathcal{T}_{val}$ and $\mathcal{T}_{test}$ are re-used from the previous task. Again, each positive subsumption in these sets is paired with $10$ negative samples, either randomly chosen or from sibling entities. Furthermore, an important factor that reflects the model’s generalisability is to examine the transfer ability across hierarchies. To this end, we extend the mixed-hop prediction task with a transfer setting where models trained on one hierarchy are tested on another.
Evaluation Metrics
For both Multi-hop Inference and Mixed-hop Prediction tasks, we utilise Precision, Recall, and F1 score (abbreviated as F-score in latter discussion) as our primary metrics of evaluation. We have opted not to include Accuracy, as preliminary testing indicated a potential bias in this metric, with a misleadingly high score resulting from the much larger volume of negative compared to positive samples. It is important to note that, although the training phase uses entity triplets, the evaluation only involves entity pairs.
4.2 Dataset Construction
We constructed the primary dataset from the noun hierarchy of WordNet (Miller, 1995) due to its comprehensive and structured representation of linguistic hierarchies. To assess the transferability and robustness across different domains, we additionally constructed datasets from ontologies that represent varied semantic granularities and domains, namely Schema.org (Guha et al., 2016), Food Ontology (FoodOn) (Dooley et al., 2018), and Disease Ontology (DOID) (Schriml et al., 2012). We retrieved WordNet from NLTK [^10] https://www.nltk.org/ and adopted pre-processing steps similar to Nickel & Kiela (2017), utilising the hypernym relations between noun synsets to construct the hierarchy. For Schema.org, FoodOn, and DOID, our pre-processing paralleled that in He et al. (2023b), transforming these ontologies into hierarchies of named entities (details in Appendix A). To accommodate the textual input requirements of LMs, we constructed an entity lexicon using the name attribute in WordNet and the rdfs:label property in the ontologies. [^11] We selected the first name (in English) if multiple names for one entity were available.
On WordNet (Noun), FoodOn, and DOID, we adopt a consistent splitting ratio for the validation and testing sets. Specifically, we allocate two separate $5\%$ portions of the indirect subsumptions $\mathcal{T}$ to form $\mathcal{T}_{val}$ and $\mathcal{T}_{test}$, respectively. Similarly, two distinct $5\%$ portions of the direct subsumptions $\mathcal{E}$ are used as $\mathcal{E}_{val}$ and $\mathcal{E}_{test}$. As Schema.org is significantly smaller than the other hierarchies and only used for transfer evaluation, we split its entire $\mathcal{E}$ and $\mathcal{T}$ sets into halves for validation and testing, respectively. Table 1 presents the extracted hierarchies’ statistics and the resulting datasets for the Multi-hop Inference and Mixed-hop Prediction tasks.
In addition to our main evaluation, we constructed a dataset from the widely-recognised biomedical ontology SNOMED CT (Stearns et al., 2001) and conducted futher evaluation. The relevant details are presented in Appendix C.
Random Negatives Hard Negatives
Model Precision Recall F-score Precision Recall F-score
NaivePrior 0.091 0.091 0.091 0.091 0.091 0.091
Multi-hop Inference (WordNet)
all-MiniLM-L6-v2 0.160 0.442 0.235 0.132 0.507 0.209
+ fine-tune 0.800 0.513 0.625 0.764 0.597 0.670
+ HiT 0.864 0.879 0.871 0.905 0.908 0.907
all-MiniLM-L12-v2 0.127 0.585 0.209 0.108 0.740 0.188
+ fine-tune 0.811 0.515 0.630 0.819 0.530 0.643
+ HiT 0.880 0.927 0.903 0.910 0.906 0.908
all-mpnet-base-v2 0.281 0.428 0.339 0.183 0.359 0.242
+ fine-tune 0.796 0.501 0.615 0.758 0.628 0.687
+ HiT 0.897 0.936 0.916 0.886 0.912 0.899
Mixed-hop Prediction (WordNet)
all-MiniLM-L6-v2 0.160 0.438 0.235 0.131 0.504 0.208
+ fine-tune 0.747 0.575 0.650 0.769 0.578 0.660
+ HiT 0.835 0.877 0.856 0.882 0.843 0.862
all-MiniLM-L12-v2 0.127 0.583 0.209 0.111 0.625 0.188
+ fine-tune 0.794 0.517 0.627 0.859 0.515 0.644
+ HiT 0.875 0.895 0.885 0.886 0.857 0.871
all-mpnet-base-v2 0.287 0.439 0.347 0.197 0.344 0.250
+ fine-tune 0.828 0.536 0.651 0.723 0.622 0.669
+ HiT 0.892 0.910 0.900 0.869 0.858 0.863
Table 2: Multi-hop Inference and Mixed-hop Prediction test results on WordNet, comparing LMs of different sizes and different settings (pre-trained, fine-tuned, and hierarchy re-trained).
4.3 Baselines
Naive Prior
We first introduce a naive baseline that utilises the prior probability of positive subsumptions in the training set for prediction. Given that each positive sample is paired with $10$ negatives, the prior probability of a positive prediction stands at $\frac{1}{11}$. Consequently, Precision, Recall, and F-score on the test set are all $\frac{1}{11}$.
Pre-trained LMs
[…]
The pre-trained LMs selected from the sentence transformer library include all-MiniLM-L6-v2, all-MiniLM-L12-v2, and all-mpnet-base-v2, each re-trained on 1B sentence pairs from a spectrum of NLP tasks. [^12] https://discuss.huggingface.co/t/train-the-best-sentence-embedding-model-ever-with-1b-training-pairs/7354 The all-MiniLM models, derived from the MiniLM distilled from BERT (Wang et al., 2020), are noted for their efficiency and performance, with all-MiniLM-L6-v2 and all-MiniLM-L12-v2 being the variants of $6$ and $12$ layers, respectively. The embedding dimension of them is $384$. The all-mpnet-base-v2 model, based on MPNet (Song et al., 2020), is selected for its top performance in the sentence transformer collection131313 https://www.sbert.net/docs/pretrained_models.html. It has an embedding dimension of $768$. While the selected models are not specifically designed for hierarchy re-training, their optimisation for clustering and semantic search tasks makes them suitable initial checkpoints.
Fine-tuned LMs
Fine-tuned LMs are used as baselines to demonstrate that despite the efficacy in various tasks, standard fine-tuning struggles to address this specific challenge. Following the BERTSubs approach outlined in Chen et al. (2023), we employ pre-trained LMs with an added linear layer for binary classification, and optimising on the Softmax loss. Chen et al. (2023) have shown that this method outperforms various structure-based embedding techniques, such as TransE (Bordes et al., 2013) and DistMult (Yang et al., 2014), and also surpasses OWL2Vec* (Chen et al., 2021), which integrates both structural and textual embeddings, in the task of predicting subsumptions.
In Appendix B, we present additional comparison with static Poincaré embeddings (StaticPoincaré) (Nickel & Kiela, 2017) on the Multi-hop Inference task. We did not consider StaticPoincaré as a primary baseline because (i) this paper focuses on addressing limitations in LMs and (ii) StaticPoincaré cannot be applied to the Mixed-hop Prediction task (and the transfer settings) due to its fixed vocabulary for entities.
4.4 Implementation
The code implementation of this work primarily depends on DeepOnto (He et al., 2023a) for processing hierarchies and constructing datasets, Geoopt (Kochurov et al., 2020) for Poincaré ball, Sentence-Transformers (Reimers & Gurevych, 2019) and Huggingface-Transformers (Wolf et al., 2020) for training and evaluation of LMs. All our experiments were conducted on a single Quadro RTX 8000 GPU.
Random Negatives Hard Negatives
Model Precision Recall F-score Precision Recall F-score
NaivePrior 0.091 0.091 0.091 0.091 0.091 0.091
Transfer Mixed-hop Prediction (WordNet $\rightarrow$ Schema.org)
all-MiniLM-L12-v2 0.312 0.524 0.391 0.248 0.494 0.330
+ fine-tune 0.391 0.433 0.411 0.597 0.248 0.351
+ HiT 0.503 0.613 0.553 0.408 0.583 0.480
Transfer Mixed-hop Prediction (WordNet $\rightarrow$ FoodOn)
all-MiniLM-L12-v2 0.135 0.656 0.224 0.099 0.833 0.176
+ finetune 0.436 0.382 0.407 0.690 0.177 0.282
+ HiT 0.690 0.463 0.554 0.741 0.385 0.507
Transfer Mixed-hop Prediction (WordNet $\rightarrow$ DOID)
all-MiniLM-L12-v2 0.342 0.451 0.389 0.159 0.455 0.235
+ fine-tune 0.585 0.621 0.603 0.868 0.179 0.297
+ HiT 0.696 0.711 0.704 0.810 0.435 0.566
Table 3: Transfer Mixed-hop Prediction test results on Schema.org, FoodOn, and DOID.
In the hierarchy re-training of our HiT models, we configured the hyperbolic clustering loss margin ($\alpha$ in Equation 3) at $5.0$ and the hyperbolic centripetal loss margin ($\beta$ in Equation 4) at $0.1$. An exception was made for all-mpnet-base-v2 with hard negatives, where $\alpha$ was adjusted to $3.0$, based on validation. The models were trained for $20$ epochs, with a training batch size of $256$, $500$ warm-up steps and an initial learning rate of $10^{-5}$, using the AdamW optimiser (Loshchilov & Hutter, 2018). Model selection was conducted after each epoch, guided by performance on the validation set.
For fine-tuning, we largely adhered to the standard settings of the Huggingface Trainer [^14] https://huggingface.co/docs/transformers/main_classes/trainer, but maintained the same training batch size as used in the hierarchy re-training. Notably, our preliminary testing revealed that fine-tuning is more prone to overfitting. Consequently, we adopted a more frequent model selection interval, performing this assessment every $500$ training steps rather than epoch-wise.
4.5 Results
The effectiveness of our hierarchy re-training approach is evident from the results of both the Multi-hop Inference and Mixed-hop Prediction tasks on WordNet (see Table 2), as well as the Transfer Mixed-hop Prediction task on Schema.org, FoodOn, and DOID for pre-trained LMs and models trained on WordNet (see Table 3). In the following, we present several pivotal findings based on these results.
Performance of HiTs
The HiT models, re-trained from LMs of various sizes, consistently outperform their pre-trained and fine-tuned counterparts across all evaluation tasks. In the Multi-hop Inference task, HiTs exhibit exceptional performance with F-scores ranging from $0.871$ to $0.916$. This indicates a strong capability in generalising from asserted to transitively inferred entity subsumptions. In the Mixed-hop Prediction task, F-scores ranging from $0.856$ to $0.900$ highlight the effectiveness of HiTs in generalising from asserted to arbitrary entity subsumptions.
For the Transfer Mixed-hop Prediction tasks, we selected all-MiniLM-L12-v2 as the pre-trained model because all-MiniLM-L12-v2+HiT attains comparable performance to all-mpnet-base-v2+HiT while it is more computationally efficient owing to a smaller parameter size. Notably, all-MiniLM-L12-v2+HiT performs better than pre-trained and fine-tuned all-MiniLM-L12-v2 on these transfer tasks by at least $0.150$ and $0.101$ in F-scores, respectively.
Limited Hierarchical Knowledge in Pre-trained LMs
For the tasks on WordNet, all-mpnet-base-v2 achieves the highest F-scores among all the pre-trained models, yet these scores (e.g., $0.347$ and $0.250$ on the Mixed-hop Prediction task with random negatives and hard negatives, respectively) are considerably lower compared to their fine-tuned (lagging by $0.304$ and $0.419$) and hierarchy re-trained (lagging by $0.553$ and $0.613$) counterparts. This disparity confirms findings from LM probing studies such as those by Hanna & Mareček (2021) and He et al. (2023b), demonstrating the limited hierarchical knowledge in pre-trained LMs.
Limited Generalisation in Fine-tuned LMs
The research by Chen et al. (2023) illustrates that fine-tuned LMs perform well on single-hop subsumptions. Our observations concur, showing that fine-tuned LMs achieve comparable performance as HiTs when assessed on just single-hop test samples. However, their effectiveness wanes when applied to arbitrary entity subsumptions. For the tasks on WordNet, fine-tuned LMs underperform HiTs by $0.194$ to $0.301$ in F-scores. In the transfer task from WordNet to Schema.org, the fine-tuned all-MiniLM-L12-v2 model only marginally outperforms its initial state, with an increase of around $0.02$ in F-scores across both negative settings.
Comparison of Random and Hard Negatives
For the tasks on WordNet, hard negative settings present greater challenges compared to random negative settings for all pre-trained LMs. This increased difficulty, however, is not as pronounced in fine-tuned LMs and HiTs. A plausible explanation is that while hard negatives pose challenges, they simultaneously act as high-quality adversarial examples, potentially leading to more robust training outcomes. In the Transfer Mixed Prediction task, hard negative settings are generally more challenging than random negative settings. For instance, in the WordNet-to-DOID transfer task, both fine-tuned and hierarchy re-trained all-MiniLM-L12-v2 models exhibit significantly higher F-scores in the random negative setting, with differences of $0.306$ and $0.138$ respectively, compared to the hard negative setting.
Case Analysis on WordNet-to-DOID Transfer
In the WordNet-to-DOID transfer task, the disparity in the “disease” category is notable: WordNet contains only $605$ entities that are descendants of “disease”, compared to over $10K$ in DOID. Despite this significant difference, HiT models effectively transfer knowledge, achieving F-scores of $0.704$ (random negatives) and $0.566$ (hard negatives).
Figure 3: Visualisation of the distribution of WordNet entity embeddings generated by HiT (re-trained from all-MiniLM-L12-v2) w.r.t. their hyperbolic norms.
4.6 Analysis of HiT Embeddings
Distribution
Figure 3 illustrates the distribution of WordNet entity embeddings generated by HiT w.r.t. their hyperbolic norms. These norms effectively capture the natural expansion of the hierarchical structure, evidenced by an exponential rise in the number of child entities. A notable observation is the sharp decline in the number of entities when hyperbolic norms exceed $23$, suggesting that few entities reside at these higher levels. Additionally, the range of entity hyperbolic norms, approximately from $8$ to $24$, indicates that a relatively small region of the high-dimensional manifold suffices to accommodate all entities in WordNet.
Case Study
In Table 4, we showcase the effectiveness of HiT using selected entities: “computer”, “pc” [^15] The full name “personal computer” is used for embedding., “fruit”, and “berry”. The table presents the hyperbolic distances between these entities’ embeddings, their individual hyperbolic norms, and their depths [^16] Depth of an entity is the minimum number of hops to the root node. For hierarchies that do not have a root node, we set up an imaginary root node when calculating the depth. in the WordNet hierarchy. We can observe that: (i) closely related entities, such as “fruit” and “berry”, are significantly nearer to each other compared to more distant pairs; (ii) more specific entities like “pc” and “berry” are positioned further from the origin of the manifold than their ancestor entities; (iii) the disparity in hyperbolic norms between “pc” and “computer” is greater compared to that between “fruit” and “berry”, reflecting the hierarchical depth where “pc” is a grandchild of “computer”, while “berry” is a direct child of “fruit”.
computer pc fruit berry
computer 0.0 5.9 22.5 24.9
pc 5.9 0.0 25.2 27.2
fruit 22.5 25.2 0.0 6.72
berry 24.9 27.2 6.72 0.0
h-norm 17.5 19.1 15.3 16.6
depth 9 11 9 10
Table 4: Hyperbolic distances between the embeddings of selected entities (“computer”, “pc”, “fruit”, “berry”), along with their individual hyperbolic norms (h-norm) and depths in WordNet.
5 Related Work
Prompt-based probing is widely used for extracting knowledge from LMs. Studies like Hanna & Mareček (2021) utilised cloze-style prompts for hypernym detection, while He et al. (2023b) approached subsumption prediction similar to Natural Language Inference. Lin & Ng (2022) examined if LMs, when correctly predicting “A is a B” and “B is a C”, can consistently infer the transitive relationship “A is a C”. These studies collectively highlight the limited capacity of pre-trained LMs in understanding hierarchical structures. Other research efforts, such as those by Liu et al. (2021) and Gosselin & Zouaq (2023), have aimed to incorporate structural semantics into LMs for entity encoding. However, these largely focus on entity equivalence or similarity, with less emphasis on hierarchical organisation.
Regarding hyperbolic embeddings, methods like the Poincaré embeddings (Nickel & Kiela, 2017) and the hyperbolic entailment cone (Ganea et al., 2018b) have shown effectiveness in representing hierarchies. However, these techniques are inherently static, limited by a fixed vocabulary size for entities, and lack inductive prediction capability – a limitation that can be mitigated by integrating them with LMs. Chen et al. (2020) ventured into this direction, applying learnable layers to project LM embeddings into hyperbolic space for syntax parsing and sentiment analysis. Our approach diverges from theirs by focusing on training LMs as general hierarchy encoders without the need for additional learnable parameters.
6 Conclusion
This paper tackles the challenge of enabling language models to interpret and encode hierarchies. We devise the hierarchy re-training approach that involves a joint optimisation on both the hyperbolic clustering and hyperbolic centripetal losses, aiming to cluster and organise entities according to their hierarchical relationships. The resulting HiT models demonstrate proficiency in simulating transitive inference and predicting subsumptions within and across hierarchies. Additionally, the analysis of HiT embeddings highlights their potential in hierarchy-oriented semantic search, which will be our future work.
Appendix A Hierarchy Construction from Ontologies
The Terminology Box (TBox) in an OWL (Web Ontology Language) [^17] https://www.w3.org/OWL/ ontology defines entity relationships using subsumption axioms ($C\sqsubseteq D$) and equivalence axioms ($C\equiv D$), where $C$ and $D$ are atomic or complex concept expressions defined in the description logic $\mathcal{SROIQ}$. In this work, we used atomic concepts as nodes of the hierarchy. We employed an ontology reasoner to deduce direct [^18] Refer to the definition of $DirectSubClassOf$ at https://owlcs.github.io/owlapi/apidocs_4/org/semanticweb/owlapi/reasoner/OWLReasoner.html. subsumptions and included these as edges in the hierarchy. This approach fully considers both subsumption and equivalence axioms for all potential edges. Considering an atomic concept $Beef$, which is defined by $Beef\equiv Meat\sqcap\exists derivesFrom.Cattle$, as an example, the reasoning will lead to an edge between $Beef$ and $Meat$. Without reasoning, $Beef$ will be misplaced under the root node, if no other subsumptions about $Beef$ are asserted in the ontology. Tools like Protégé [^19] https://protege.stanford.edu/ also demonstrate similar considerations in presenting ontology concept taxonomies.
It is also important to note the variability in naming schemes across different ontologies, which sometimes necessitates pre-processing of entity names. He et al. (2023b) provided detailed pre-processing steps for Schema.org, FoodOn, and DOID. In this study, we used the pre-processed versions of FoodOn and DOID and applied the same pre-processing methodology to the latest version of Schema.org.
Appendix B Results for Static Poincaré Embeddings
Random Negatives Hard Negatives
Model Precision Recall F-score Precision Recall F-score
Multi-hop Inference (WordNet)
StaticPoincaré 0.862 0.866 0.864 0.797 0.867 0.830
all-MiniLM-L12-v2 0.127 0.585 0.209 0.108 0.740 0.188
+ fine-tune 0.811 0.515 0.630 0.819 0.530 0.643
+ HiT 0.880 0.927 0.903 0.910 0.906 0.908
Table 5: Multi-hop Inference test results on WordNet, comparing StaticPoincaré and the HiT model re-trained from all-MiniLM-L12-v2.
Table 5 provides additional comparison to the static Poincaré embeddings (StaticPoincaré) in the Multi-hop Inference task on WordNet. Although StaticPoincaré surpasses the pre-trained and fine-tuned all-MiniLM-L12-v2 in terms of F-scores, it still falls short of all-MiniLM-L12-v2+HiT, by margins of $0.039$ and $0.105$ in the random and hard negative settings, respectively. This comparison further highlights the effectiveness of integrating hyperbolic geometry with language models.
A crucial advantage of HiT over StaticPoincaré is its capability for inductive prediction, enabling its application in tasks like Mixed-hop Prediction and transfer learning scenarios, where StaticPoincaré is not applicable due to its static vocabulary for entities.
Notably, the random negative setting in the Multi-hop Inference task follows the evaluation framework in Ganea et al. (2018b). Our implementation of StaticPoincaré not only outperforms their baseline models but also shows better results than their hyperbolic entailment cone model, even though the latter was trained with an extra $10\%$ of multi-hop edges.
Appendix C Results on SNOMED CT
In the main body of this paper, we primarily focused on models trained on the WordNet (Noun). This section broadens our study to encompass models trained on SNOMED CT (Stearns et al., 2001), a structured, comprehensive, and widely-used vocabulary for electronic health records. Using the tool provided by the SNOMED CT team, [^20] https://github.com/IHTSDO/snomed-owl-toolkit we converted the latest version of SNOMED CT (released in December 2023) into the OWL ontology format. We then constructed the hierarchy according to the procedure detailed in Appendix A.
[…]
Source #Entity #DirectSub #IndirectSub #Dataset (Train/Val/Test)
SNOMED 364,352 420,193 2,775,696 mixed: 4,160K///1,758K///1,758K
Table 6: Statistics of SNOMED-CT, including the numbers of entities (#Entity), direct subsumptions (#DirectSub), indirect subsumptions (#IndirectSub), and the dataset splittings (#Dataset) for Multi-hop Inference and Mixed-hop Prediction tasks.
In Table 6, we present relevant statistics of the SNOMED CT hierarchy and its corresponding dataset, which was constructed using the method outlined in Section 4.2.
Random Negatives Hard Negatives
Model Precision Recall F-score Precision Recall F-score
MajorityPrior 0.091 0.091 0.091 0.091 0.091 0.091
Mixed-hop Prediction (SNOMED)
all-MiniLM-L12-v2 0.224 0.443 0.297 0.145 0.398 0.213
+ fine-tune 0.919 0.859 0.888 0.894 0.635 0.743
+ HiT 0.941 0.967 0.954 0.905 0.894 0.899
Transfer Mixed-hop Prediction (SNOMED $\rightarrow$ Schema.org)
all-MiniLM-L12-v2 0.312 0.524 0.391 0.248 0.494 0.330
+ fine-tune 0.198 0.864 0.322 0.431 0.288 0.345
+ HiT 0.432 0.580 0.495 0.274 0.608 0.378
Transfer Mixed-hop Prediction (SNOMED $\rightarrow$ FoodOn)
all-MiniLM-L12-v2 0.135 0.656 0.224 0.099 0.833 0.176
+ fine-tune 0.378 0.540 0.445 0.638 0.371 0.469
+ HiT 0.700 0.500 0.583 0.594 0.442 0.506
Transfer Mixed-hop Prediction (SNOMED $\rightarrow$ DOID)
all-MiniLM-L12-v2 0.342 0.451 0.389 0.159 0.455 0.235
+ fine-tune 0.547 0.912 0.684 0.831 0.795 0.812
+ HiT 0.836 0.864 0.850 0.739 0.748 0.744
Table 7: Mixed-hop Prediction test results on SNOMED, comparing LMs of different sizes and different settings (pre-trained, fine-tuned, and hierarchy re-trained), and Transfer Mixed-hop Prediction results on Schema.org, FoodOn, and DOID.
Table 7 details the Mixed-hop Prediction results on SNOMED CT, along with the Transfer Mixed-hop Prediction results on Schema.org, FoodOn, and DOID for pre-trained LMs and models trained on SNOMED CT. Building on the demonstrated effectiveness of HiT in simulating transitive inference from the main body of this paper, we present only the results for inductive subsumption prediction. The transfer evaluation incorporates the same set of hierarchies as those used in the evaluation on WordNet (Noun), facilitating a meaningful comparison of model performance across different training hierarchies.
In the Mixed-hop Prediction task on SNOMED CT, all models – pre-trained, fine-tuned, and hierarchy re-trained – outperform their counterparts on WordNet (Noun) in terms of F-scores. This enhanced performance is likely attributed to SNOMED CT’s more specific entity naming, reducing homonyms and thus ambiguity. In the transfer results, models trained on SNOMED CT exhibit notably better F-scores on DOID, which aligns with expectations given DOID’s focus on diseases and SNOMED CT’s comprehensive biomedical scope. Conversely, their performance on Schema.org, a common-sense ontology, is comparatively worse. Notably, the fine-tuned all-MiniLM-L12-v2 performs even worse than its pre-trained version on Schema.org, suggesting an overfitting to biomedical domain knowledge in SNOMED CT.